OpenStudy (anonymous):

Divide the complex number of 8+i over 2-i

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

\[\frac{ 8+i }{2-i },rationalise \it\]

OpenStudy (anonymous):

Is it 16i+3i-i2

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

\[\frac{ 8+i }{2-i }*\frac{ 2+i }{ 2+i }=\frac{ 16+8i+2i+i ^{2} }{2^{2}-i ^{2}}=\frac{ 16+10i-1 }{ 4+1 }=\frac{ 15+10i }{ 5 }=3+2i\]

OpenStudy (anonymous):

if there is any problem ask me.

OpenStudy (anonymous):

Where the 8i come from

OpenStudy (anonymous):

multiply 2+i by 8

OpenStudy (anonymous):

So the answe is 16 over what?

OpenStudy (anonymous):

answer is 3+2i

OpenStudy (anonymous):

i am leaving now .it is already 11.14PM

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."